# Fluid Antenna Array Enhanced Over-the-Air Computation

### v1

@@ Abstract
Over-the-air computation (AirComp) has emerged as a promising technology for fast wireless data aggregation by harnessing the superposition property of wireless multiple-access channels. This paper investigates a fluid antenna (FA) array-enhanced AirComp system, employing the new degrees of freedom achieved by antenna movements. Specifically, we jointly optimize the transceiver design and antenna position vector (APV) to minimize the mean squared error (MSE) between target and estimated function values. To tackle the resulting highly non-convex problem, we adopt an alternating optimization technique to decompose it into three subproblems. These subproblems are then iteratively solved until convergence, leading to a locally optimal solution. Numerical results show that FA arrays with the proposed transceiver and APV design significantly outperform the traditional fixed-position antenna arrays in terms of MSE.
@@ I Introduction
This paper addresses this gap by investigating an FA array-enhanced AirComp system, wherein an AP equipped with an FA array aggregates data concurrently transmitted from multiple users via AirComp. Our objective is to minimize the mean squared error (MSE) between target and estimated function values by jointly optimizing the transceiver design and antenna position vector (APV). To handle the resulting highly non-convex problem, we adopt an alternating optimization (AO) technique to decompose it into three subproblems, corresponding to 1) the transmit equalization coefficients of users, 2) the AP decoding vector, and 3) the APV. The first two subproblems are convex quadratically constrained quadratic programs (QCQPs) and can be efficiently solved using off-the-shelf solvers. However, optimizing the APV remains a non-convex problem, for which we employ the primal-dual interior point (PDIP) method. By iteratively solving these subproblems until convergence, we achieve a locally optimal solution for the original problem. Numerical results demonstrate that FA arrays with the proposed transceiver and APV design significantly outperform traditional FPAs in terms of MSE.
@@ II-B Problem Formulation
where (7b) accounts for the maximum transmission power constraint for each user, (7c) guarantees that the FAs are moved within the feasible region $[0,L]$, and (7d) ensures that the distance between two adjacent FAs is no less than $L_{0}$ to avoid antenna coupling.
@@ III Transceiver and APV Design
In the sequel, we adopt the AO technique to resolve the coupling between $\bm{b}$, $\bm{m}$, and $\bm{x}$ in (II-B), and optimize one variable at a time with others being fixed.
@@
It is observed that $b_{1},\cdots,b_{K}$ in (III) are decoupled and can be decomposed into $K$ subproblems. The subproblem associated with $b_{k}$, $\forall k\in\mathcal{K}$, is given by
@@
Firstly, the Lagrangian associated with (III) can be expressed as follows:
@@
Before solving (III), we rewrite (17a) into a more tractable form. Firstly, we expand $\sum\nolimits_{k=1}^{K}\left|\bm{m}^{H}\bm{a}(\bm{x},\theta_{k})\alpha_{k}b_{k}-1\right|^{2}$ as follows:
@@
$=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}|w_{k,n}w_{k,l}|e^{j[\phi_{k}(x_{n}-x_{l})-(\angle w_{k,n}-\angle w_{k,l})]}$
$=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}|w_{k,n}w_{k,l}|\cos[q_{k}(x_{n},x_{l})],$ (19)
@@
Algorithm 1 converges to a locally optimal point of problem (II-B) after several iterations.
@@
The proof is analogous to that for [Algorithm 1, 9], and thus we omit it for brevity.
@@
1:Given $b_{k}=\sqrt{P_{k}}$, $\forall k\in\mathcal{K}$, and $x_{n}=\frac{Ln}{N+1}$, $\forall n\in\{1,\cdots,N\}$;
@@ IV Numerical Results
In this section, we present numerical results demonstrating the effectiveness of introducing FA arrays in enhancing AirComp performance. We set $L_{0}=0.5\lambda$, $L=8\lambda$, $P_{k}=P_{0}$, $\forall k\in\mathcal{K}$, and $\text{SNR}\triangleq P_{0}/\sigma^{2}$. Moreover, the following benchmark schemes are considered for comparison.
  • Fixed-Position Antenna (FPA): The $N$ FAs are fixed and uniformly distributed in the feasible region $[0,L]$, i.e., the APV is fixed to $\bm{x}=\left[\frac{L}{N+1},\cdots,\frac{NL}{N+1}\right]^{T}$. The optimal $\bm{b}$ and $\bm{m}$ are obtained by recursively running (12) and (16) until convergence.
  • SCA: Referring to [11], the non-convex optimization problem (III) for $\bm{x}$ can also be solved by using the successive convex approximation (SCA) technique, while the optimization procedures for $\bm{b}$ and $\bm{m}$ remain unchanged, i.e., (12) and (16), respectively. For more details, refer to Appendix B.
In Fig. 2, we show the MSE performance across iterations for the aforementioned schemes under different SNR levels. From this figure, we immediately observe that our proposed scheme significantly outperforms the FPA scheme for both SNR values. In contrast, the SCA scheme only achieves marginal improvement compared to the FPA scheme, particularly when $\text{SNR}=0$ dB, which can be attributed to the relaxations used in constructing convex surrogate functions for (17a). Additionally, we observe from this figure that all three schemes exhibit rapid convergence.
Figure 2: MSE versus the number of iterations under different SNRs, where $N=8$ and $K=10$.
Figure 3: MSE versus the number of iterations for different numbers of FAs $N$, where $\text{SNR}=-10$ dB, and $K=10$.
Fig. 3 depicts the MSE performance for the three aforementioned schemes across different numbers of FAs, i.e., $N$. As expected, the MSE performance consistently decreases across all three schemes as $N$ increases from $4$ to $12$. It is also observed that our proposed scheme showcases a significant advantage over the FPA scheme at both $N=4$ and $N=12$, again highlighting the benefits of optimizing the positions of FAs.
Figure 4: MSE versus the number of iterations for different numbers of users, where $N=8$ and $\text{SNR}=-10$ dB.
In Fig. 4, the MSE performance for the three aforementioned schemes across different user numbers, i.e., $K$, is presented. Unsurprisingly, the MSE performance consistently rises across all three schemes as $K$ increases. Notably, we also observe the consistent outperformance of the proposed scheme compared to the two benchmark schemes, with the performance gap widening as $K$ increases.
@@ V Conclusions
In this paper, we studied an FA array-enhanced AirComp system. Compared to traditional FPAs, FAs provide new DoFs through antenna movements, thus offering the potential for improved performance. As such, we jointly optimized the transceiver design and APV to minimize the MSE between target and estimated function values. Numerical results demonstrated that FA arrays with our proposed transceiver and APV design significantly outperformed traditional FPAs in terms of MSE.
@@
where $\bm{e}_{n}$ denotes the $n$-th column of $\bm{I}_{N}$, $\forall n\in\{1,\cdots,N\}$. Given $g(\bm{x})\triangleq\sum\nolimits_{k=1}^{K}[F_{k}(\bm{x})-G_{k}(\bm{x})]$ and $f_{i}(\bm{x})$, $\forall i=1,\cdots,N+1$, we rewrite (III) as follows:
@@
The Lagrangian associated with (A) is given by
@@
$\begin{bmatrix}\nabla^{2}g(\bm{x})+\sum\nolimits_{i=1}^{N+1}\nu_{i}\nabla^{2}f_{i}(\bm{x})&D\bm{f}(\bm{x})^{T}\\[4.30554pt] -{\rm diag}(\bm{\nu})D\bm{f}(\bm{x})&-{\rm diag}(\bm{f}(\bm{x}))\end{bmatrix}\begin{bmatrix}\Delta\bm{x}\\[4.30554pt] \Delta\bm{\nu}\end{bmatrix}$ (29)
$=-\begin{bmatrix}\bm{r}_{\rm dual}(\bm{x},\bm{\nu})\\[4.30554pt] \bm{r}_{\rm cent}(\bm{x},\bm{\nu})\end{bmatrix}.~{}~{}~{}~{}~{}~{}~{}$ (32)
@@
Algorithm 2 Pseudo-Code for Solving (A) with the PDIP Method
@@
4:     Compute $(\Delta\bm{x}_{\rm pd},\Delta\bm{\nu}_{\rm pd})$ via solving (29);
@@
$\varphi^{i}_{k,n,l}=|w_{k,n}w_{k,l}|[\sin(q_{k}(x_{n}^{i},x_{l}^{i}))\phi_{k}+\phi^{2}_{k}(x^{i}_{n}-x^{i}_{l})],$
$C^{i}_{k}=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}|w_{k,n}w_{k,l}|[0.5\phi^{2}_{k}(x^{i}_{n}-x^{i}_{l})^{2}$
@@
where $\bar{w}_{k,0}=\sum\nolimits_{n=1}^{N}|w_{k,n}|$, $\bar{\bm{w}}_{k}=|\bm{w}_{k}|$, and $q_{k}(x_{n}^{i},x_{l}^{i})=\phi_{k}(x^{i}_{n}-x^{i}_{l})-(\angle w_{k,n}-\angle w_{k,l})$.
@@
$[\tilde{\bm{v}}^{i}_{k}]_{n}=|w_{k,n}|[\phi^{2}_{k}x^{i}_{n}-\sin(\phi_{k}x_{n}^{i}-\angle w_{k,n})\phi_{k}],$
$\tilde{C}^{i}_{k}=\sum\limits_{n=1}^{N}|w_{k,n}|[\cos(\phi_{k}x_{n}^{i}-\angle w_{k,n})$
@@
where $\bm{A}=\sum\nolimits_{k=1}^{K}(\bm{A}_{k}+\tilde{\bm{A}}_{k})$, $\bm{v}^{i}=\sum\nolimits_{k=1}^{K}(\bm{v}^{i}_{k}+2\tilde{\bm{v}}^{i}_{k})$, and $C^{i}=\sum\nolimits_{k=1}^{K}(C^{i}_{k}-2\tilde{C}^{i}_{k}+1)$. Since (7c) and (7d) are linear constraints and $\bm{A}$ is a positive semi-definite matrix, therefore (B) is a convex optimization problem and can be efficiently solved using CVX [15].

### v2

@@ Abstract
Over-the-air computation (AirComp) has emerged as a promising technology for fast wireless data aggregation by harnessing the superposition property of wireless multiple-access channels. This paper investigates a fluid antenna (FA) array-enhanced AirComp system, employing the new degrees of freedom introduced by antenna movements. Specifically, we jointly optimize the transceiver design and antenna position vector (APV) to minimize the mean squared error (MSE) between target and estimated function values. To tackle the resulting highly non-convex problem, we adopt an alternating optimization technique to decompose it into three subproblems. These subproblems are then iteratively solved until convergence, leading to a locally optimal solution. Numerical results show that FA arrays with the proposed transceiver and APV design significantly outperform the traditional fixed-position antenna arrays in terms of MSE.
@@ I Introduction
This paper addresses this gap by investigating an FA array-enhanced AirComp system, wherein an AP equipped with an FA array aggregates data concurrently transmitted from multiple users via AirComp. Our main contributions are summarized as follows:
1) We seek to minimize the mean squared error (MSE) between target and estimated function values by jointly optimizing the transceiver design and antenna position vector (APV). To handle the resulting highly non-convex problem, we adopt an alternating optimization (AO) technique to decompose it into three subproblems, corresponding to $a)$ the transmit equalization coefficients of users, $b)$ the AP decoding vector, and $c)$ the APV.
2) The first two subproblems are convex quadratically constrained quadratic programs (QCQPs) and can be efficiently solved using off-the-shelf solvers. However, optimizing the APV remains a non-convex problem, for which we employ the primal-dual interior point (PDIP) method. By iteratively solving these subproblems until convergence, we achieve a locally optimal solution for the original problem.
3) Numerical results demonstrate that FA arrays with the proposed transceiver and APV design significantly outperform traditional FPAs in terms of MSE.
@@ II-B Problem Formulation
where (7b) accounts for the power constraint for each user, with $P_{k}$ denoting the maximum transmission power of user $k$, $\forall k\in\mathcal{K}$, (7c) guarantees that the FAs are moved within the feasible region $[0,L]$, and (7d) ensures that the distance between two adjacent FAs is no less than $L_{0}$ to avoid antenna coupling.
@@ III Transceiver and APV Design
In the sequel, we adopt the AO technique to resolve the coupling between $\bm{b}$, $\bm{m}$, and $\bm{x}$ in (7), and optimize one variable at a time with others being fixed.
@@
It is observed that $b_{1},\cdots,b_{K}$ in (8) are decoupled and can be decomposed into $K$ subproblems. The subproblem associated with $b_{k}$, $\forall k\in\mathcal{K}$, is given by
@@
Firstly, the Lagrangian associated with (9) can be expressed as follows:
@@
Before solving (17), we rewrite (17a) into a more tractable form. Firstly, we expand $\sum\nolimits_{k=1}^{K}\left|\bm{m}^{H}\bm{a}(\bm{x},\theta_{k})\alpha_{k}b_{k}-1\right|^{2}$ as follows:
@@
$=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}|w_{k,n}||w_{k,l}|e^{j[\phi_{k}(x_{n}-x_{l})-(\angle w_{k,n}-\angle w_{k,l})]}$
$=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}|w_{k,n}||w_{k,l}|\cos[q_{k}(x_{n},x_{l})],$ (19)
@@
Algorithm 1 converges to a locally optimal point of problem (7) after several iterations.
@@
To prove the convergence of Algorithm 1, we introduce superscript $t$ as the iteration index, e.g., $\bm{m}^{t}$ represents the decoding vector at the end of the $t$-th iteration round. Then, Algorithm 1 converges as
$\mathbb{MSE}(\bm{m}^{t},\bm{b}^{t},\bm{x}^{t})$ $\overset{(a)}{\geq}\mathbb{MSE}(\bm{m}^{t+1},\bm{b}^{t},\bm{x}^{t})$
$\overset{(b)}{\geq}\mathbb{MSE}(\bm{m}^{t+1},\bm{b}^{t+1},\bm{x}^{t})$
$\overset{(c)}{\geq}\mathbb{MSE}(\bm{m}^{t+1},\bm{b}^{t+1},\bm{x}^{t+1}),$
where $(a)$, $(b)$, and $(c)$ follow since the updates of $\bm{m}$, $\bm{b}$, and $\bm{\Phi}$ are the optimal (or locally optimal) solutions to (14), (8), and (17), respectively. Since $\mathbb{MSE}(\bm{m}^{t},\bm{b}^{t},\bm{x}^{t})$ is monotonically non-increasing in each iteration and its value is lower bounded by zero, we prove that Algorithm 1 will converge to a local optimum of problem (7) after several iterations. ∎
@@
1:Given $b_{k}=\sqrt{P_{k}}$, $\forall k\in\mathcal{K}$, and $x_{n}=\frac{Ln}{N-1}$, $\forall n\in\{0,\cdots,N-1\}$;
@@ IV Numerical Results
This section presents numerical results demonstrating the effectiveness of introducing FA arrays in enhancing AirComp performance. We set $L_{0}=0.5\lambda$, $L=N\lambda$, $P_{k}=P_{0}$, $\forall k\in\mathcal{K}$, and $\text{SNR}\triangleq P_{0}/\sigma^{2}$. Moreover, the following three benchmark schemes are considered for comparison.
  • Fixed-Position Antenna Array (FPA): The $N$ FAs are fixed and uniformly distributed in the feasible region $[0,L]$, i.e., the APV is fixed to $\bm{x}=\frac{L}{N-1}\times\left[0,1,\cdots,N-1\right]^{T}$. The optimal $\bm{b}$ and $\bm{m}$ are obtained by recursively running (12) and (16) until convergence.
  • SCA: We optimize $\bm{x}$ by solving (17) via the successive convex approximation (SCA) technique [11], while the optimization procedures for $\bm{b}$ and $\bm{m}$ remain unchanged. For more details, refer to Appendix B.
  • PGD: Problem (17) for $\bm{x}$ can also be solved by using the projected gradient descent (PGD) algorithm [14], while the optimization procedures for $\bm{b}$ and $\bm{m}$ remain unchanged.
(a) MSE versus the number of iterations, where $\text{SNR}=-10$ dB, $N=10$, and $K=100$.
(b) MSE versus SNR, where $N=10$, and $K=10$.
(c) MSE versus the number of FAs $N$, where $\text{SNR}=-10$ dB, and $K=10$.
(d) MSE versus the number of users, where $\text{SNR}=-10$ dB and $N=10$.
Figure 2: MSE performance versus the number of iterations, SNR, the number of FAs, and the number of users
To show the convergence of Algorithm 1, we present the MSE performance across iterations in Fig. 2(a). From this figure, we observe that Algorithm 1 converges after a few dozens of iterations. In contrast, the SCA scheme takes $100$ iterations to converge. Additionally, we observe that Algorithm 1 is superior to the three benchmark schemes in terms of the MSE performance.
In Fig. 2(b), we show the MSE performance across different SNR levels. From this figure, it is observed that our proposed scheme (i.e., Algorithm 1) outperforms the FPA scheme for all SNR values, though the performance gap becomes narrower as the SNR increases.
Fig. 2(c) depicts the MSE performance across different numbers of FAs, i.e., $N$. As expected, the MSE performance consistently decreases as $N$ increases from $5$ to $25$. It is also observed that our proposed scheme showcases a significant advantage over the FPA scheme, again highlighting the benefits of optimizing the positions of FAs.
Fig. 2(d) depicts the MSE performance across different user numbers, i.e., $K$. Unsurprisingly, the MSE performance consistently rises as $K$ increases. Notably, we also observe the consistent outperformance of the proposed scheme compared to the SCA and FPA schemes, with the performance gap widening as $K$ increases.
@@ V Conclusions
This paper studied an FA array-enhanced AirComp system. Compared to traditional FPAs, FAs provide new DoFs through antenna movements, thus offering the potential for improved performance. As such, we jointly optimized the transceiver design and APV to minimize the MSE between target and estimated function values. Numerical results demonstrated that FA arrays with our proposed transceiver and APV design significantly outperformed traditional FPAs in terms of MSE.
Thanks to the benefits of reducing AirComp MSE, FA arrays can also be used to improve the performance of wireless federated learning systems that rely on AirComp, which is an interesting topic for future work.
@@
where $\bm{e}_{n}$ denotes the $n$-th column of $\bm{I}_{N}$, $\forall n\in\{1,\cdots,N\}$. Given $g(\bm{x})\triangleq\sum\nolimits_{k=1}^{K}[F_{k}(\bm{x})-G_{k}(\bm{x})]$ and $f_{i}(\bm{x})$, $\forall i=1,\cdots,N+1$, we rewrite (17) as follows:
@@
The Lagrangian associated with (21) is given by
@@
$\begin{bmatrix}\nabla^{2}g(\bm{x})+\sum\nolimits_{i=1}^{N+1}\nu_{i}\nabla^{2}f_{i}(\bm{x})&D\bm{f}(\bm{x})^{T}\\[4.30554pt] -{\rm diag}(\bm{\nu})D\bm{f}(\bm{x})&-{\rm diag}(\bm{f}(\bm{x}))\end{bmatrix}\begin{bmatrix}\Delta\bm{x}\\[4.30554pt] \Delta\bm{\nu}\end{bmatrix}~{}$
$=-\begin{bmatrix}\bm{r}_{\rm dual}(\bm{x},\bm{\nu})\\[4.30554pt] \bm{r}_{\rm cent}(\bm{x},\bm{\nu})\end{bmatrix}.$ (25)
@@
Algorithm 2 Pseudo-Code for Solving (21) with the PDIP Method
@@
4:     Compute $(\Delta\bm{x}_{\rm pd},\Delta\bm{\nu}_{\rm pd})$ via solving (A);
@@
$\varphi^{i}_{k,n,l}={\bar{w}}_{k,n}{\bar{w}}_{k,l}[\sin(q_{k}(x_{n}^{i},x_{l}^{i}))\phi_{k}+\phi^{2}_{k}(x^{i}_{n}-x^{i}_{l})],$
$C^{i}_{k}=\sum\limits_{n=1}^{N}\sum\limits_{l=1}^{N}{\bar{w}}_{k,n}{\bar{w}}_{k,l}[0.5\phi^{2}_{k}(x^{i}_{n}-x^{i}_{l})^{2}$
@@
where ${\bar{w}}_{k,n}=|w_{k,n}|$, $\forall k,n$, $\bar{\bm{w}}_{k}=[{\bar{w}}_{k,1},\cdots,{\bar{w}}_{k,N}]^{T}$, $\bar{w}_{k,0}={\bar{w}}_{k,1}+\cdots+{\bar{w}}_{k,N}$, and $q_{k}(x_{n}^{i},x_{l}^{i})=\phi_{k}(x^{i}_{n}-x^{i}_{l})-(\angle w_{k,n}-\angle w_{k,l})$.
@@
$[\tilde{\bm{v}}^{i}_{k}]_{n}=\bar{w}_{k,n}[\phi^{2}_{k}x^{i}_{n}-\sin(\phi_{k}x_{n}^{i}-\angle w_{k,n})\phi_{k}],$
$\tilde{C}^{i}_{k}=\sum\limits_{n=1}^{N}\bar{w}_{k,n}[\cos(\phi_{k}x_{n}^{i}-\angle w_{k,n})$
@@
where $\bm{A}=\sum\nolimits_{k=1}^{K}(\bm{A}_{k}+\tilde{\bm{A}}_{k})$, $\bm{v}^{i}=\sum\nolimits_{k=1}^{K}(\bm{v}^{i}_{k}+2\tilde{\bm{v}}^{i}_{k})$, and $C^{i}=\sum\nolimits_{k=1}^{K}(C^{i}_{k}-2\tilde{C}^{i}_{k}+1)$. Since (7c) and (7d) are linear constraints and $\bm{A}$ is a positive semi-definite matrix, therefore (31) is a convex optimization problem and can be efficiently solved using CVX [16].